# Finch members

### v3

@@ Abstract
We introduce a finance & accounting benchmark (Finch) for evaluating AI agents on real-world, enterprise-grade professional workflows—interleaving data entry, structuring, formatting, web search, cross-file retrieval, calculation, modeling, validation, translation, visualization, and reporting. Finch is sourced from authentic enterprise workspaces at Enron (15,000 spreadsheet files and 500,000 emails from 150 employees) and other financial institutions, preserving in-the-wild messiness across multimodal artifacts (text, tables, formulas, charts, code, and images) and spanning diverse domains such as budgeting, trading, and asset management.
We propose a workflow construction process that combines LLM-assisted discovery with expert annotation: (1) LLM-assisted, expert-verified derivation of workflows from real-world email threads and version histories of spreadsheet files, and (2) meticulous expert annotation for workflows, requiring over 700 hours of domain-expert effort. This yields 172 composite workflows with 384 tasks, involving 1,710 spreadsheets with 27 million cells, along with PDFs and other artifacts, capturing the intrinsically messy, long-horizon, knowledge-intensive, and collaborative nature of real-world enterprise work.
We conduct both human and automated evaluations of frontier AI systems including GPT 5.1, Claude Sonnet 4.5, Gemini 3 Pro, Grok 4, and Qwen 3 Max. GPT 5.1 Pro spends 16.8 minutes per workflow yet passes only 38.4% of workflows, while Claude Sonnet 4.5 passes just 25.0%. Comprehensive case studies further surface the challenges that real-world enterprise workflows pose for AI agents.
@@ 1 Introduction
Frontier AI systems are increasingly transforming professional workspaces. AI-assisted tools like ChatGPT [openai_agent], Claude [claude_excel], Gemini [gemini_workspace], and Copilot [microsoft_copilot] are now embedded in daily enterprise workflows—helping professionals draft documents, explore data, manipulate spreadsheets, and generate reports. These tools are particularly impactful in finance and accounting (F&A), a high-stakes, knowledge- and labor-intensive domain critical to every organization.
However, real-world F&A work is inherently messy, with substantial contextual complexity: artifacts are interconnected across heterogeneous spreadsheets, PDFs, and other artifacts, evolving through multiple versions with collaborative edits [klimt2004enron]; spreadsheets contain large, complex structures [spreadsheetllm2024] with cross-sheet references, intricate layouts, inconsistent formatting, cryptic terms, erroneous formulas, and multimodal artifacts such as charts, images, and code. It is also long-horizon [gdpval2025]: workflows demand multi-step reasoning spanning data entry, editing, retrieval, calculation, modeling, validation, reporting, and more.
@@
  • In-the-wild enterprise sourcing: Finch is built around authentic enterprise spreadsheets, emails, and PDFs from real-world enterprise workspaces—primarily Enron [enrondata_pst] (about 15,000 spreadsheet files and 500,000 emails from 150 executives and employees) and EUSES [fisher2005euses] (about 450 financial spreadsheet files from various sources), along with securities and asset management firms, global organizations such as World Bank [worldbank_idr_2024], and Canadian and British governments [financecanada_fiscalref_2025, hmt_pesa_2023]. Documents are large, cross-referenced, and messy—containing rich multimodal artifacts such as tables, formulas, charts, pivots, and images.
@@
We compile 172 meticulously annotated, enterprise-grade workflows built on 1,710 spreadsheets, along with PDFs and other artifacts, collectively capturing the compositional, messy, knowledge-intensive, and collaborative nature of real work. Each workflow spans one or more interdependent tasks—data entry, editing, retrieval, calculation, modeling, validation, translation, visualization, and reporting—mirroring how professionals actually work on artifact manipulation and creation.
@@
We evaluate a spectrum of frontier AI systems—including Claude Sonnet 4.5, GPT 5.1, Gemini 3, Grok 4, and Qwen 3—using both expert evaluation and a novel automated evaluation pipeline that closely aligns with expert judgments. Our experiments reveal that even the frontier agents pass fewer than 40% of workflows (GPT 5.1 Pro spends 16.8 minutes per workflow on average), highlighting the substantial challenges that Finch poses for AI agents. Comprehensive case analyses further surface concrete challenges that real-world enterprise workflows pose for AI agents.
Figure 3: Illustration of an end-to-end predictive modeling workflow typically performed by financial analysts. The workflow involves multiple steps, including web search, data import, cross-sheet and cross-file retrieval, calculation and financial modeling, visualization, and report generation. More illustrative examples for data characteristics in Finch are presented in Appendix E.
@@ 2.1 Dataset Construction
We propose a novel workflow construction pipeline grounded in the real collaborative context of emails and versioned artifacts, as illustrated in Figure 4 (a–c). First, we induce workflows from enterprise email threads and versioned documents, where collaborators naturally describe, discuss, and track work as part of their daily routines. Second, we derive workflows by analyzing changes across versioned spreadsheets, surfacing the actual data transformations and analysis steps that professionals performed. Third, we leverage high-quality spreadsheets and reports: we design workflows, author task instructions, and revise these spreadsheets and reports so that they serve as the input files and reference solutions.
@@ 2.1.3 Workflow Sourced from Final Deliverable Spreadsheets and Reports
Finally, we curate workflows from high-quality spreadsheets and reports drawn from the Enron and EUSES corpora, various investment and securities companies, international organizations, and national governments (e.g., the World Bank and the Canadian and British governments). Domain experts author realistic workflow instructions and construct input and reference files based on final deliverable artifacts. For example, a valuation model from an investment firm can be turned into a financial modeling task; a World Bank report can be used to define a data-driven summarization and visualization task; and bilingual reports from the Canadian government can be used to construct translation and consistency-checking tasks. We additionally leverage labeled samples from existing datasets: we adapt 10 financial cases from WideSearch [wong2025widesearch] into web-search-centric workflows and extend them into multi-step calculation and visualization pipelines, and we leverage 3 examples from DABStep [egg2025dabstep] to construct multi-source question answering workflows.
@@ 2.1.4 Quality Control
Given the high complexity of each workflow, rigorous quality control is essential. We perform inter-annotator validation on all workflows and use the ChatGPT 5.1 Pro and Claude-Sonnet-4.5 product-side agents as secondary checkers: given the instruction and the input and reference files, the model is asked to judge whether the reference output is consistent with the instruction and whether any obvious steps are missing. LLM-based judgments are used to flag potential defects for human review. Together, these quality-control procedures yield a collection of workflows whose NL instructions, input files, and reference outputs are well aligned.
@@ 2.2 Dataset Characteristics
Finch comprises 172 meticulously annotated, enterprise-grade workflows that collectively capture the compositional, messy, multimodal, and collaborative nature of real finance and accounting work. Across these workflows, the corpus contains 1,710 spreadsheets (956 distinct spreadsheets) together with 17 PDFs, 12 images, 3 Word documents, and additional files such as JSON, CSV, and Markdown. This mixture reflects how real analysts coordinate over heterogeneous artifacts rather than clean, single-table inputs.
@@ 2.2.1 Task Compositionality
Finch is explicitly designed around composite workflows rather than isolated tasks. As shown in Figure 5, only 37 workflows (21.5%) are single-task; the remaining 135 (78.5%) involve multiple tasks. Importantly, each “task” itself typically requires substantial multi-turn reasoning: for example, web search often entails many rounds of LLM calls to discover, filter, and verify evidence; cross-sheet retrieval requires iterative calls to read and locate key information across multiple sheets; and calculation usually spans many formulas distributed over different rows and columns.
These tasks are not independent subtasks, but are interleaved around shared spreadsheets and files. A typical workflow may begin with structuring or importing raw data, proceed to cross-sheet or cross-file retrieval, and then culminate in calculations, modeling, or reporting. The distribution in Figure 5 shows that most workflows weave multiple tasks.
Figure 6: Distribution of the number of sheets and cells per workflow.
@@ 2.3.2 LLM-as-Judge Evaluation
This LLM-as-judge framework not only automates large-scale evaluation but also surfaces subtle spreadsheet errors (such as formulas silently replaced with static values) that are difficult to catch with GUI-based human inspection alone. In Section 3.2, we report the consistency between human and automated evaluations and show that the LLM-as-judge scores closely track human judgments.
@@ 3.1.1 Product-side Agents
We evaluate two frontier product-side agents: (i) ChatGPT using the GPT 5.1 model in Pro mode, and (ii) Claude using the Sonnet 4.5 model in thinking mode. We focus on these two systems rather than alternatives such as Gemini or Grok because they natively support returning downloadable files (e.g., spreadsheets) as outputs, rather than emitting code or markdown-formatted tables that are not intuitive for human evaluation. For both agents, we enable their external web browsing, but disable using historical chats so that each workflow is evaluated independently and without cross-run leakage. Since model updates are frequent and manual evaluation is very time-consuming, we used the latest model from our final round of experiments and did not consider subsequent updates.
@@ 3.1.2 API-based Models
We evaluate five frontier LLMs via API interfaces (Table 1). We adopt SpreadsheetBench [ma2024spreadsheetbench] as the baseline framework because it provides a principled code-generation paradigm for spreadsheet-centric tasks, treating executable code as the model’s action space and enabling direct manipulation of spreadsheets through standard libraries. This design naturally aligns with Finch, where workflows require complex spreadsheet operations, formula reasoning, and cross-sheet dependencies that cannot be reliably handled by text-only outputs.
@@
GPT 5.1 [gpt5openai] OpenAI 400K 128K
Claude Sonnet 4.5 [claudesonnet45] Anthropic 1M 64K
Grok 4 [grok4] xAI 256K 256K
Qwen 3 Max [qwen3technicalreport] Alibaba 256K 32.8K
Gemini 3 Pro Preview [gemini3pro] Google 1.05M 65.5K
@@
SpreadsheetBench produces text tables without preserving cell addresses, data types, or formulas. However, these details are essential for tasks in Finch. We extend SpreadsheetLLM [spreadsheetllm2024] encoding and introduce a semantic-rich tuple encoding that preserves full structural and semantic fidelity. Each sheet begins with its name and the corresponding data bounding box (e.g. ## Sheet: [name] (A1:Z100)). We then serialize the bounded region using a Markdown-based format. Each cell is encoded as a tuple (Address, Value, Type, Formula), where Address denotes the cell reference (e.g. A3), Type indicates the data type (T = Text, I = Integer, F = Float, D = Date, B = Boolean), and Formula records the cell formula (e.g., =SUM(A1:A10)->100).
@@ 3.2 Experimental Results
Product-side agents (ChatGPT 5.1 Pro vs. Claude Sonnet 4.5).
As shown in Figure 2 and Table 3, ChatGPT 5.1 Pro achieves the best overall pass rates on Finch. Their advantage largely comes from rich interactive affordances: they can iteratively inspect spreadsheets, revise intermediate states, and recover from partial errors over many tool calls. However, it solves fewer than 40% of the Finch workflows, suggesting that real-world finance and accounting work remains far from “solved” even for frontier agents. The detailed analysis in Figure 8 further highlights that long-horizon composition is a key bottleneck: when a workflow contains more than two tasks, the pass rate drops sharply—GPT 5.1 Pro decreases from 44.3% (workflows with $\leq 2$ tasks) to 23.5% (workflows with >2>2 tasks), and Claude Sonnet 4.5 decreases from 30.3% to 11.8%. This indicates that error accumulation across steps and missing intermediate affordances disproportionately hurt multi-step execution.
@@
We further analyze how GPT 5.1 web completion time scales with the number of tasks in a workflow (Table 2). The longest individual workflow run takes roughly 60 minutes to complete—but fails, highlighting how challenging workflows can be for current agents. On average, single-task workflows take 13.1 minutes, while workflows with two, three, and four tasks require 17.4, 18.7, and 21.3 minutes, respectively, reflecting the increased compositional complexity of multi-task settings. Interestingly, five-task workflows show a lower average time (13.6 minutes), most of which involve web search and almost all of which result in failure.

### v4

@@ Abstract
We introduce FinWorkBench (a.k.a. Finch) for evaluating agents on real-world, enterprise-grade finance and accounting workflows—interleaving data entry, structuring, formatting, web search, cross-file retrieval, calculation, modeling, validation, translation, visualization, and reporting. Finch is sourced from authentic enterprise workspaces from Enron (15,000 files and 500,000 emails) and across various financial institutions, covering the period 2000–2025, preserving the in-the-wild messiness of multimodal artifacts such as tables and charts, across diverse domains including budgeting, trading, and asset management.
We propose a workflow construction process that combines LLM-assisted mining of workflows from authentic enterprise environments with expert annotation: (1) LLM-assisted, expert-verified derivation of workflows from real-world email threads and version histories of spreadsheet files, and (2) meticulous annotation for workflows, requiring over 700 hours of expert effort. This yields 172 composite workflows with 384 tasks, involving 1,710 spreadsheets with 27 million cells, along with PDFs and other artifacts, capturing the intrinsically messy, long-horizon, knowledge-intensive, and collaborative nature of enterprise work.
We conduct both human and automated evaluations of frontier AI systems, including GPT 5.1, Claude Sonnet/Opus 4.5, Gemini 3 Pro, Grok 4, and Qwen 3 Max. GPT 5.1 Pro spends an average of 16.8 minutes per workflow yet passes only 38.4% of workflows. Comprehensive case studies further surface the challenges that real-world enterprise workflows pose for AI agents.
@@ 1 Introduction
Frontier AI systems are increasingly transforming professional workspaces. AI-assisted tools like Claude [1], OpenClaw [42], ChatGPT [41], Gemini [22], and Copilot [39] are now embedded in daily enterprise workflows—helping professionals draft documents, explore data, manipulate spreadsheets, and generate reports. These tools are particularly impactful in finance and accounting (F&A), a high-stakes, knowledge- and labor-intensive domain critical to every organization.
However, real-world F&A work is inherently messy, with substantial contextual complexity: artifacts are interconnected across heterogeneous spreadsheets, PDFs, and other artifacts, evolving through multiple versions with collaborative edits [28]; spreadsheets contain large, complex structures [17] with cross-sheet references, intricate layouts, inconsistent formatting, cryptic terms, erroneous formulas, and multimodal artifacts such as charts, images, and code. It is also long-horizon [43]: workflows demand multi-step reasoning spanning data entry, editing, retrieval, calculation, modeling, validation, reporting, and more.
@@
  • In-the-wild enterprise sourcing: Finch is built around authentic enterprise spreadsheets, emails, and PDFs from real-world enterprise workspaces—primarily Enron [19] (about 15,000 spreadsheet files and 500,000 emails from 150 executives and employees) and EUSES [20] (about 450 financial spreadsheet files from various sources), along with financial institutions, global organizations, and government agencies, including the International Debt Report 2025, World Bank 2024 reports, Adidas 2024 financial statement, Public Accounts of Canada 2024 [4, 11, 24], etc. Documents are large, cross-referenced, and messy—containing rich multimodal artifacts such as tables, formulas, charts, pivots, and images.
@@
We compile 172 meticulously annotated, enterprise-grade workflows built on 1,710 spreadsheets, along with PDFs and other artifacts, collectively capturing the compositional, messy, knowledge-intensive, and collaborative nature of real work. Each workflow spans one or more interdependent tasks—data entry, editing, retrieval, calculation, modeling, validation, translation, visualization, and reporting—mirroring how professionals actually work on artifact manipulation and creation. Beyond their compositional task structure, Finch workflows demand substantial multi-step agent interaction: a case study on 20 representative workflows reveals that each requires a median of 16 tool calls (range: 6–107).
@@
Evaluating such workflows poses nontrivial challenges, because Finch tasks usually involve complex and large spreadsheets and require assessing both structural correctness and semantic fidelity beyond exact matching. To address this, we provide both expert human evaluation and a scalable LLM-as-judge pipeline. Human evaluation serves as the gold standard, judging whether a workflow has been satisfactorily completed end-to-end. In parallel, we introduce an efficient multimodal LLM-as-judge evaluation method that compares inputs, model outputs, and reference artifacts using structured diffs, compact snapshots, and multimodal renderings, enabling reliable assessment at scale.
We evaluate a spectrum of frontier AI systems—including Claude Sonnet/Opus 4.5, GPT 5.1, Gemini 3, Grok 4, and Qwen 3—using both expert evaluation and our automated pipeline. Even the strongest frontier agents pass fewer than 50% of Finch workflows: GPT 5.1 Pro spends an average of 16.8 minutes per workflow yet achieves only 38.4% under human evaluation.
Case analyses further reveal how each dimension of real-world complexity contributes to failure. Long-horizon composition is a key bottleneck: GPT 5.1 Pro drops from 44.3% on workflows with $\leq$2 tasks to 23.5% on those with >>2 tasks, as errors accumulate across steps. The messy, irregular structure of enterprise spreadsheets leads to frequent data retrieval errors, and agents often struggle to reconstruct latent knowledge-intensive business logic encoded in spreadsheet formulas. Multimodal artifacts further amplify difficulty: on the 20 workflows involving PDFs, images, or Word documents, GPT 5.1 Pro’s pass rate drops to 35.0%, below its overall rate. Taken together, these results suggest that it is the combination of messy inputs, long-horizon dependencies, domain knowledge, and heterogeneous formats—rather than any single factor—that drives the sharp performance degradation on real enterprise workflows.
Figure 3: Illustration of an end-to-end predictive modeling workflow typically performed by financial analysts. The workflow involves multiple steps, including web search, data import, cross-sheet and cross-file retrieval, calculation and financial modeling, visualization, and report generation. More illustrative examples for data characteristics in Finch are presented in Appendix F.
@@ 2.1 Dataset Construction
Echoing the existentialist dictum that existence precedes essence [44], we argue that professional work should be observed in real-world enterprise environment before it is formally defined. Motivated by this perspective, we propose a novel workflow construction pipeline grounded in the authentic collaborative context of emails, versioned spreadsheets, and final deliverable spreadsheets and reports, as illustrated in Figure 4 (a–c). First, we induce workflows from enterprise email threads and versioned documents, where collaborators naturally describe, discuss, and track work as part of their daily routines. Second, we derive workflows by analyzing changes across versioned spreadsheets, surfacing the actual data transformations and analysis steps that professionals performed. Third, we leverage high-quality spreadsheets and reports: we design workflows, author task instructions, and revise these spreadsheets and reports so that they serve as the input files and reference solutions.
@@ 2.1.3 Workflow Sourced from Final Deliverable Spreadsheets and Reports
Finally, we curate workflows from high-quality spreadsheets and reports drawn from the Enron and EUSES corpora, various investment and securities companies, international organizations, and national governments (e.g., the World Bank and the Canadian and British governments). Domain experts author realistic workflow instructions and construct input and reference files based on final deliverable artifacts. For example, a valuation model from an investment firm can be turned into a financial modeling task; a World Bank report can be used to define a data-driven summarization and visualization task; and bilingual reports from the Canadian government can be used to construct translation and consistency-checking tasks. We additionally leverage labeled samples from existing datasets: we adapt 10 financial cases from WideSearch [49] into web-search-centric workflows and extend them into multi-step calculation and visualization pipelines, and we leverage 3 examples from DABStep [18] to construct multi-source question answering workflows.
@@ 2.1.4 Quality Control
Given the high complexity of each workflow, we adopt a rigorous multi-stage quality control process. All workflows are annotated and reviewed by a team of five experts: two annotators with interdisciplinary backgrounds in finance and computer science, and three annotators with computer science backgrounds. Among them, three annotators (including one female) have over nine years of industry experience, and two are outstanding Ph.D. or master’s students.
Annotators are instructed to skip workflows that are similar to existing ones to maximize diversity. 20% of our annotators are full-time reviewers who do not create any original annotations, and 80% act as full-time workflow creators (producing the initial annotations) and also serve as part-time reviewers. Each workflow is proposed by one person and typically reviewed by two others, including a dedicated full-time reviewer who reviews all queries and also inspects LLM-generated review comments to surface potential annotation issues. Importantly, Quality Control (QC) is iterative rather than one-pass: approximately 40% of workflows undergo at least one round of revision; 20+ workflows went through three or more QC rounds, and we held regular group discussions to resolve challenging cases and align standards. In addition, ChatGPT 5.1 Pro and Claude Sonnet 4.5 are used as secondary checkers to flag potential issues, which are always verified by human experts. Together, these procedures required over 700 hours of expert annotation.
@@ 2.2 Dataset Characteristics
Finch comprises 172 meticulously annotated, enterprise-grade workflows that collectively capture the compositional, messy, multimodal, and collaborative nature of real finance and accounting work. Across these workflows, the corpus contains 1,710 spreadsheets (956 distinct ones in 301 Excel files) together with 17 PDFs, 12 images, 3 Word documents, as well as JSON, CSV, and Markdown. This mixture reflects how real analysts coordinate over heterogeneous artifacts rather than clean, single-table inputs.
@@ 2.2.1 Task Compositionality
Finch is explicitly designed around composite workflows rather than isolated tasks. As shown in Figure 5, only 37 workflows (21.5%) are single-task; the remaining 135 (78.5%) involve multiple tasks. These tasks are not independent subtasks, but are interleaved around shared spreadsheets and files. A typical workflow may begin with structuring or importing raw data, proceed to cross-sheet or cross-file retrieval, and then culminate in calculations, modeling, or reporting. The distribution in Figure 5 shows that most workflows weave multiple tasks.
Importantly, each “task” itself typically requires substantial multi-turn reasoning: for example, web search often entails many rounds of LLM calls to discover, filter, and verify evidence; cross-sheet retrieval requires iterative calls to read and locate key information across multiple sheets; and calculation usually spans many formulas distributed over different rows and columns.
To further characterize the practical complexity of Finch workflows, we conduct a case study on tool-call overhead using Claude Coworker (Opus 4.6) on 20 representative tasks (Appendix D). Excluding two extremely tool-intensive web-search workflows that require 71 and 107 tool calls respectively, the remaining tasks range from 6 to 25 tool calls (median 14). Notably, tool-call overhead does not scale linearly with the number of tasks; intrinsic task complexity—such as the depth of business-logic reasoning and cross-sheet dependencies—is a stronger driver. These findings suggest that even single- or two-task workflows can demand substantial multi-step agent interaction, reinforcing that Finch captures meaningful operational complexity beyond what task counts alone convey.
@@ 2.2.2 Messiness
Figure 6: Distribution of the number of sheets and cells per workflow.
@@ 2.3.2 LLM-as-Judge Evaluation
This LLM-as-judge framework not only automates large-scale evaluation but also surfaces subtle spreadsheet errors (such as formulas silently replaced with static values) that are difficult to catch with GUI-based human inspection alone. In Section 3.2, we report the consistency between human and automated evaluations and show that the LLM-as-judge scores closely track human judgments. Looking forward, agentic evaluation is quite promising for future work for flexible workflow evaluation.
@@ 3.1.1 Product-side Agents
We evaluate two frontier product-side agents: (i) ChatGPT using the GPT 5.1 model in Pro mode, and (ii) Claude using the Sonnet/Opus 4.5 models in thinking mode. We focus on these two systems rather than alternatives such as Gemini or Grok because they natively support returning downloadable files (e.g., spreadsheets) as outputs, rather than emitting code or markdown-formatted tables that are not intuitive for human evaluation. For both agents, we enable their external web browsing, but disable using historical chats so that each workflow is evaluated independently and without cross-run leakage. Since model updates are frequent and manual evaluation is very time-consuming, we used the latest model from our final round of experiments and did not consider subsequent updates.
@@ 3.1.2 API-based Models
We evaluate five frontier LLMs via API interfaces (Table 1). We adopt SpreadsheetBench [37] as the baseline framework because it provides a principled code-generation paradigm for spreadsheet-centric tasks, treating executable code as the model’s action space and enabling direct manipulation of spreadsheets through standard libraries. This design naturally aligns with Finch, where workflows require complex spreadsheet operations, formula reasoning, and cross-sheet dependencies that cannot be reliably handled by text-only outputs.
@@
GPT 5.1 [40] OpenAI 400K 128K
Claude Sonnet 4.5 [3] Anthropic 1M 64K
Grok 4 [51] xAI 256K 256K
Qwen 3 Max [55] Alibaba 256K 32.8K
Gemini 3 Pro Preview [21] Google 1.05M 65.5K
@@
SpreadsheetBench produces text tables without preserving cell addresses, data types, or formulas. However, these details are essential for tasks in Finch. We extend SpreadsheetLLM [17] encoding and introduce a semantic-rich tuple encoding that preserves full structural and semantic fidelity. Each sheet begins with its name and the corresponding data bounding box (e.g. ## Sheet: [name] (A1:Z100)). We then serialize the bounded region using a Markdown-based format. Each cell is encoded as a tuple (Address, Value, Type, Formula), where Address denotes the cell reference (e.g. A3), Type indicates the data type (T = Text, I = Integer, F = Float, D = Date, B = Boolean), and Formula records the cell formula (e.g., =SUM(A1:A10)->100).
@@ 3.2 Experimental Results
Product-side agents (ChatGPT 5.1 Pro vs. Claude Opus/Sonnet 4.5).
As shown in Figure 2 and Table 5, ChatGPT 5.1 Pro and Claude Opus 4.5 achieve the strongest overall pass rates on Finch. Their advantage largely comes from rich interactive affordances: they can iteratively inspect spreadsheets, revise intermediate states, and recover from partial errors over many tool calls. However, they solve fewer than 50% of the Finch workflows, suggesting that real-world finance and accounting work remains far from “solved” even for frontier agents. The detailed analysis in Figure 8 further highlights that long-horizon composition is a key bottleneck: when a workflow contains more than two tasks, the pass rate drops sharply—GPT 5.1 Pro decreases from 44.3% (workflows with $\leq 2$ tasks) to 23.5% (workflows with >2>2 tasks), and Claude Sonnet 4.5 decreases from 30.3% to 11.8%. This indicates that error accumulation across steps and missing intermediate affordances disproportionately hurt multi-step execution.
@@
We further analyze how GPT 5.1 web completion time scales with the number of tasks in a workflow (Table 2). The longest individual workflow run takes roughly 60 minutes to complete—but fails, highlighting how challenging workflows can be for current agents. On average, single-task workflows take 13.1 minutes, while workflows with two and three tasks require 17.4 and 18.7 minutes, respectively, reflecting the increased compositional complexity of multi-task settings. Interestingly, workflows with 4 or more tasks show a lower average time (17.4 minutes) than workflows with 3 tasks, most of which involve web search and almost all of which result in failure.
@@
In addition, we examine performance on a subset of 20 workflows that involve non-spreadsheet artifacts (e.g., PDFs, Word documents, and images). As shown in Table 3, GPT 5.1 Pro achieves a 35.0% pass rate on these workflows (7/20), lower than its overall rate of 38.4%, while Claude Sonnet 4.5 achieves 25.0% (5/20), matching its overall rate. This suggests that handling heterogeneous document types introduces additional challenges.
Pass Rate Overall Other file types
GPT-5.1 Pro (Product) 38.4% 35.0%
Claude Sonnet 4.5 (Product) 25.0% 25.0%
Table 3: Pass rates on workflows involving non-spreadsheet artifacts (e.g., PDFs, Word documents, and images) under human evaluation.
Beyond binary pass/fail, our LLM-as-judge framework provides rubric-level scores along four dimensions: completeness, correctness, over-edit avoidance, and readability. Table 4 reports these scores for the two Claude product-side agents. Correctness is the primary bottleneck: Sonnet 4.5 achieves 55.8% completeness but only 36.0% correctness. These scores are produced by GPT-5-mini using structured output, where a single call assigns all four scores jointly; this may introduce inter-rubric confusion. More robust agentic evaluation methods are a worthwhile direction for future work.
Rubric (%) Sonnet 4.5 Opus 4.5
Completeness 55.8 69.8
Correctness 36.0 45.3
Over-edit Avoidance 55.8 64.0
Readability 65.1 74.4
Table 4: Rubric-level scores under automated evaluation.